For the function defined by f(x)=2-x^2, 0 is < and = to x, I need to find a formula for f^-1(x).
This is real frustrating...

Answer by praseenakos@yahoo.com(507)

QUESTION:

f(x)=2-x^2, 0 is < and = to x, I need to find a formula for f^-1(x).

ANSWER:


f(x)=2-x^2


For our convenience, lets take,


y = 2-x^2


Rewrite the given equation as follows.



Add x^2 to both sides of the equation,



==> y + x^2 = 2-x^2 + x^2



==> x^2 + y = 2



Subtract y from both sides of the equation,




==> x^2 + y - y = 2 - y



==> x^2 = 2 - y




Take square root on both sides of the expression,




==> x = square root of ( 2 - y)




We can write this expressions(after replacing y by x ) as,



f-1(x) = sqrt (2-x)



Which is the required inverse function.




Here we can see that this inverse function is not defined when the value
of x is greater than 2 because then 2-x becomes a negative number and sqrt of
a negative number cannot be determined.



So we can write, inverse of the given function is sqrt of (2-x) and x is less
than or equal to 2.




Hope you Understood.



Regards.


praseenakos@yahoo.co.in